# UU-Mamba: Uncertainty-aware U-Mamba for Cardiovascular Segmentation

### v1

@@ 3.1 Mamba Block and U-Mamba Network
The U-Mamba network is designed to improve the accuracy of medical image segmentation and improve global context comprehension by combining the assets of the Mamba block [41] and U-Net [37, 38]. The Mamba block, which is specifically engineered for Selective Structured State Space Sequence Models (S6), is particularly well-suited for medical imaging duties due to its exceptional ability to manage long-range dependencies and sequential information.
@@
State Space Models (SSM) [49] describe systems in terms of their internal states and observations over time, thereby facilitating effective sequence modeling through these underlying states. The fundamental form is denoted as follows: $\mathbf{x}_{t}$ is the input state vector, $\mathbf{u}_{t}$ is the control input, $\mathbf{w}_{t}$ is the process noise, $\mathbf{A}$ is the state transition matrix, and $\mathbf{B}$ is the control input matrix.
@@
For observation $\mathbf{y}_{t}$, calculated using the observation matrix $\mathbf{C}$, feedthrough matrix $\mathbf{D}$, and observation noise $\mathbf{v}_{t}$, the formula is:
@@
The integration of S6 into the Mamba block is especially crucial for sequential medical image processing tasks, such as cardiac MRI segmentation, which require the capture of temporal dynamics and structure [41]. The method, on the other hand, is exclusively concerned with per-image segmentation, which involves the application of the state transition and observation matrices ($\mathbf{A}$, $\mathbf{C}$, etc.) to individual images. Each image is treated independently.
U-Mamba capitalizes on Mamba’s linear scaling advantage to improve CNNs’ capacity to simulate long-range dependencies, all while circumventing the high computational costs associated with self-attention mechanisms employed in Transformers [50] such as ViT [42] and SwinTransformer [51]. The U-Mamba block, which is comprised of two sequential residual blocks followed by a Mamba block, is depicted in Figure 2.
@@
An encoder with these blocks is included in the complete U-Mamba network architecture to capture both local features and long-range dependencies, while a decoder composed of residual blocks and transposed convolutions is used to recover detailed local information and resolution. Skip connections are used to connect hierarchical features from the encoder to the decoder. The final decoder output is processed through a 1$\times$1$\times$1 convolutional layer and a Softmax layer to generate the final segmentation probability map.
3.2 Uncertainty-aware Loss
Introducing uncertainty into loss functions entails allocating weights to distinct components of the loss according to the estimated uncertainty for each data point [52, 53]. This method allows the model to concentrate on learning from more dependable instances while simultaneously reducing the impact of potentially erroneous or ambiguous data. Kendall and Gal introduced the concept of adjusting loss functions by utilizing homoscedastic and heteroscedastic uncertainty [54]. Heteroscedastic uncertainty fluctuates between instances, while homoscedastic uncertainty remains constant across all data points. The model can improve its resilience and precision by focusing on confident predictions and reducing the impact of equivocal ones by adapting its learning process to capitalize on these uncertainties. This optimization enhances overall performance and improves the training process across diverse datasets [55, 56, 57].
@@ 3.2 Uncertainty-aware Loss
    Dice loss [15]: This region-based metric emphasizes the overlap between predicted and ground truth areas, ensuring accurate preservation of shape and boundary details in segmented regions.
@@
    Cross-Entropy (CE) loss [16]: This distribution-based loss ensures precise categorization of individual pixels, thereby improving classification accuracy.
@@
    Focal loss [17]: This pixel-level loss addresses class imbalance by assigning greater importance to challenging instances, enhancing the model’s ability to manage complex scenarios [58, 59, 60].
@@
3.3 Sharpness-Aware Minimization Optimization
To improve the U-Mamba model’s generalizaiton in segmenting cardiovascular images, such as those in the ACDC [2], ImageCAS [9], and Aorta dataset [10, 11], our methodology employs Sharpness-Aware Minimization (SAM) optimization [13]. The model’s generalizability is enhanced by the flattening of the loss landscape, implemented by SAM optimization. Despite the fact that conventional optimization techniques are designed to identify the lowest points in the loss landscape, these points are frequently precipitous, which results in inadequate generalization to new data. SAM, on the other hand, identifies gentler minima—regions in the parameter space where the model’s performance remains consistent and is less susceptible to perturbations.
@@ 3.3 Sharpness-Aware Minimization Optimization
By applying these two stages iteratively, SAM directs the optimizer toward parameter configurations that are resilient to perturbations, thereby improving generalization and overall performance. The model is able to identify flattened minima in the loss landscape as a result of this approach, which results in improved generalization [13].
Figure 3: Segmentation results for various methods on sample images from the ACDC dataset [2]. The Mean Squared Error (MSE) between the output segmentation and the ground truth is shown for each method.
Figure 4: Segmentation results for various methods on sample images from the Aorta dataset [10, 11]. The Mean Squared Error (MSE) between the output segmentation and the ground truth is shown for each method.
Figure 5: Segmentation results for various methods on sample images from the ImageCAS dataset [9]. The Mean Squared Error (MSE) between the output segmentation and the ground truth is shown for each method.
4 Experiments
4.1 Experimental Settings
4.1.1 The ACDC Dataset
The Automated Cardiac Diagnosis Challenge (ACDC) dataset [2] is a widely recognized benchmark in medical image analysis, particularly for cardiac MRI segmentation. With a total of 300 images and 2,978 slices, this dataset comprises MRI scans from 150 patients, each divided into numerous slices. normal subjects, myocardial infarction, dilated cardiomyopathy, hypertrophic cardiomyopathy, and abnormal right ventricle are the five distinct categories in which the patients are evenly distributed. Each group is distinguished by specific cardiac pathologies.
@@ 4.1.1 The ACDC Dataset
4.1.2 The Aorta Dataset
The Aorta dataset [10, 11] is a meticulously annotated compilation of 50 Computed Tomography Angiography (CTA) images that enables the multi-class segmentation of the aorta and its branches. The axial dimensions of these images range from $389\times 389$ pixels to $516\times 516$ pixels, with an average size of $450\times 450$ pixels. The dataset is guaranteed to be consistent in measurement, as each image maintains an isotropic voxel resolution of 1mm x 1mm x 1mm. The average number of axial segments per scan is 695, with image numbers ranging from 578 to 801. This dataset is essential for the development and testing of sophisticated algorithms that are designed to accurately segment the complex vascular structures within the aorta and its branches, thereby establishing a strong foundation for research in medical image analysis and machine learning.
4.1.3 The ImageCAS Dataset
The ImageCAS dataset [9] focuses on the segmentation of coronary arteries using CTA images. This dataset contains approximately 1,000 3D CTA images, making it considerably larger than existing public datasets in this domain, which are crucial for diagnosing and assessing coronary artery disease. The dataset is particularly challenging due to the small size and complex branching patterns of the coronary arteries, as well as the motion artifacts introduced by cardiac and respiratory movements. Ground truth annotations include detailed segmentations of the coronary arteries, providing a comprehensive framework for evaluating the performance of segmentation algorithms in detecting and delineating these critical structures.
@@ 4.1.3 The ImageCAS Dataset
4.1.4 Evaluation Metrics
We employ the Dice Similarity Coefficient (DSC) as our primary metric for evaluating segmentation performance, as per the evaluation protocol outlined in [26]. The DSC assesses the overlap between the predicted segmentation and the ground truth mask, thereby providing a reliable indication of the model’s accuracy in delineating cardiac structures. The DSC is defined in Eq. (4).
@@ 4.1.4 Evaluation Metrics
We also incorporate the Normalized Surface Dice (NSD) metric that recommend in [26]. The NSD measures the average distance between the predicted segmentation surface and the ground truth surface, normalized by the ground truth surface area. This metric is particularly useful for assessing the spatial accuracy of the segmentation, especially in clinical scenarios where precise boundary delineation is critical. The NSD is defined as follows:
@@ 4.1.5 Implementation Details
We conducted the experiments using the PyTorch framework and two NVIDIA A100 Tensor Core GPUs for training. During training, for the ACDC dataset [2], we used a patch size of [20, 256, 224] and a batch size of 4, with the number of pooling operations per axis configured to [2, 5, 5]. In the case of the ImageCAS dataset [9], a patch size of [96, 160, 160] and a batch size of 2 were selected, with pooling operations per axis set to [4, 5, 5]. For the Aorta dataset [10, 11], the patch size was [176, 112, 112], also with a batch size of 2, and the pooling operations per axis were set to [4, 4, 4]. The network configuration comprises 6 stages. An initial learning rate of $5\times 10^{-3}$ was utilized, and the training proceeded for 500 epochs. In the SAM optimization, the hyperparameter $\rho$ controlling the perturbation in Eq. (9) was set to 0.05. The focusing parameter $\gamma$ in the focal loss in Eq. (7) was set to 2. The parameter $M$ of the uncertainty-aware loss in Eq. (8) was set to 3, incorporating Dice loss, Cross-Entropy loss, and Focal loss.
Method Average RV $\uparrow$ Myo $\uparrow$ LV $\uparrow$ TransUNet [61] 89.71 88.86 84.53 95.73 Swin-Unet [62] 90.00 88.55 85.62 95.83 nnUNet [37] 91.61 90.24 89.24 95.36 nnFormer [63] 92.06 90.94 89.58 95.65 U-Mamba [12] 92.22 91.83 90.22 94.54 UU-Mamba (Ours) 92.79 92.41 90.90 95.04
@@
Table 1: Performance comparison of our UU-Mamba with leading medical image segmentation methods on the ACDC dataset [2] for the three anatomical regions—the right ventricle (RV), left ventricle (LV), and myocardium (Myo). The evaluation metric is DSC (%).
Method Avg. DSC % $\uparrow$ RV DSC % $\uparrow$ Myo DSC % $\uparrow$ LV DSC % $\uparrow$ CE loss 92.263 91.81 90.31 94.67 Uncertainty-aware loss (CE, Dice, Focal) 92.602 92.36 90.51 94.94 UU-Mamba model (Ours) 92.787 92.41 90.90 95.04
@@
Table 2: The U-Mamba backbone was employed to conduct an ablation study of ACDC dataset [2] for the three anatomical regions—the right ventricle (RV), left ventricle (LV), and myocardium (Myo). The study included the following configurations: (1) only the Cross-Entropy (CE) loss, (2) the uncertainty-aware loss without the SAM optimizer, and (3) the proposed UU-Mamba model (uncertainty-aware loss + SAM optimizer). DSC (%) serves as the evaluation metric.
Method Avg. DSC % $\uparrow$ Avg. NSD % $\uparrow$ Avg. MSE % $\downarrow$ CE loss 73.761 92.141 0.0960 Uncertainty-aware loss (CE, Dice, Focal) 75.053 91.747 0.0966 UU-Mamba model (Ours) 77.084 93.847 0.0906
Table 3: Ablation study of Aorta dataset [10, 11] on various configurations with U-Mamba backbone: (1) using only the Cross-Entropy (CE) loss, (2) using the uncertainty-aware Loss without the SAM optimizer, and (3) the proposed UU-Mamba model (uncertainty-aware loss + SAM optimizer). The evaluation metric is average DSC (%), NSD (%), and MSE.
@@
Method Avg. DSC % $\uparrow$ Avg. NSD % $\uparrow$ Avg. MSE % $\downarrow$ CE loss 79.496 87.490 0.0006784 Uncertainty-aware loss (CE, Dice, Focal) 81.146 88.045 0.0006105 Uncertainty-aware loss + SAM (Ours) 81.9983 88.771 0.0005903
@@
Table 4: Ablation study of ImageCAS dataset [9] on various configurations with U-Mamba backbone: (1) using only the Cross-Entropy (CE) loss, (2) using the uncertainty-aware Loss without the SAM optimizer, and (3) the proposed UU-Mamba model. The evaluation metric is average DSC (%), NSD (%), and MSE.
4.2 Experimental Results
@@
We conduct a comparison of UU-Mamba with five of the state-of-the-art segmentation models—TransUNet [61], Swin-Unet [62], nnUNet [37], nnFormer [63], and U-Mamba [12]—on the ACDC dataset [2]. Transformer-based networks are TransUNet and Swin-Unet, while nnUNet and nnFormer employ CNN-based architectures. U-Mamba is a hybrid architecture that combines components from both Transformer-based and CNN-based networks.
@@
Using the Dice Similarity Coefficient (DSC) as the evaluation metric, we conducted a quantitative evaluation of UU-Mamba against these five 3D heart segmentation models on the ACDC dataset [2]. The segmentation results for the compared algorithms are depicted in Figure 3 on a few images from the ACDC dataset. Table 1 provides the average DSC scores across all regions, as well as the DSC scores for each model in three cardiac regions: the right ventricle (RV), myocardium (Myo), and left ventricle (LV). The scores that demonstrate the greatest performance are indicated in italics.
@@
Our UU-Mamba model surpasses all other methods, attaining the highest overall performance with an average DSC of 92.79%. The DSC scores for each region are as follows: 92.41% for RV, 90.90% for Myo, and 95.04% for LV for each region. Our model’s flexibility and efficacy are demonstrated by its exceptional ability to accurately segment the right ventricle and myocardium. Despite a minor decrease in DSC for the left ventricle in comparison to other models, this is counterbalanced by the highest overall average DSC and the superior scores in other regions, as illustrated in Table 1.
@@
In contrast, TransUNet attains an average DSC of 89.71%, with a relatively lower score for Myo. The average DSC of Swin-Unet is 90.00%, with the maximum DSC for LV and lower performance for RV. The nnUNet model achieves an average DSC of 91.61%, with significant enhancements in Myo segmentation. Providing robust performance in all regions, particularly the myocardium, the nnFormer model obtains an average DSC of 92.06%. U-Mamba’s average DSC of 92.22% indicates substantial improvements in the RV and Myo regions.
@@
The superior performance of our UU-Mamba model in comparison to the other existing models is emphasized by this quantitative evaluation. Our method exhibits the potential to improve the accuracy of 3D heart segmentation in medical imaging, as it has the highest average DSC and notably strong segmentation in the right ventricle and myocardium. The effectiveness of our approach is underscored by the enhancements it achieves over models such as U-Mamba, nnFormer, and nnUNet. This approach employs sophisticated loss functions and optimization techniques to capitalize on both global and local features.
DSC (%) NSD (%) Labels CE loss Uncertainty- aware loss UUMamba CE loss Uncertainty- aware loss UUMamba Zone 0 89.317 87.803 87.604 76.749 70.280 72.291 Innominate 75.151 77.830 78.126 81.251 84.259 84.768 Zone 1 67.936 65.432 65.491 86.647 84.379 85.001 Left Common Carotid 78.001 77.547 78.689 92.852 93.539 93.453 Zone 2 75.043 75.514 73.749 91.645 91.627 90.585 Left Subclavian Artery 83.260 83.578 84.725 99.124 98.742 98.612 Zone 3 74.306 73.513 73.157 94.146 92.939 93.038 Zone 4 79.590 84.233 84.701 89.008 91.987 93.001 Zone 5 88.860 89.725 89.780 95.571 96.879 97.271 Zone 6 71.166 71.119 73.552 97.692 94.694 98.381 Celiac Artery 67.016 66.311 68.691 97.270 96.138 99.158 Zone 7 68.116 71.036 74.067 99.404 99.060 99.622 SMA 69.002 70.117 71.200 88.940 87.300 88.905 Zone 8 77.029 79.027 78.831 100.000 99.576 100.000 Right Renal Artery 74.873 72.898 73.501 98.032 96.949 97.015 Left Renal Artery 67.907 70.187 71.890 96.650 98.323 97.640 Zone 9 90.805 90.118 91.008 99.877 99.636 99.637 Right Common Iliac Artery 78.244 79.078 86.141 95.438 96.317 98.783 Left Common Iliac Artery 75.914 79.623 86.457 97.638 96.370 99.864 Right Internal Iliac Artery 59.568 65.540 66.409 85.795 82.728 88.424 Left Internal Iliac Artery 67.034 64.154 67.791 96.124 90.865 99.069 Right External Iliac Artery 59.216 61.344 72.469 77.545 77.270 87.139 Left External Iliac Artery 59.148 70.506 74.895 81.853 90.306 96.825
@@
Table 5: Ablation study of comparing the DSC and NSD values for various anatomical zones and arteries using three methods: (1) only the Cross-Entropy (CE) loss, (2) the uncertainty-aware Loss without the SAM optimizer, and (3) the proposed UU-Mamba model. The best values for each region are highlighted in bold. The table demonstrates the effectiveness of the UU-Mamba model in achieving higher segmentation accuracy across most regions.
@@
Figure 6: Comparison of loss landscapes for models with and without SAM (Sharpness-Aware Minimization) optimization across three datasets: ACDC [2], ImageCAS [9], and Aorta [10, 11]. Each image represents the 3D loss landscape of a model trained on one of the datasets, illustrating the effect of SAM optimization on the model’s ability to find flatter minima. The models without SAM exhibit sharper, more erratic loss contours, indicating less stable convergence, whereas the models with SAM show smoother, reflecting enhanced generalization capabilities. This visualization highlights the impact of SAM optimization in improving model robustness and training stability across diverse data conditions.
4.3 Ablation Study

### v2

@@ III-B Network Architecture
The U-Mamba network is designed to improve the accuracy of medical image segmentation and improve global context comprehension by combining the assets of the Mamba block [gu2023mamba] and U-Net [Isensee2021nnUNet, ronneberger2015u]. The Mamba block, which is specifically engineered for Selective Structured State Space Sequence Models (S6), is particularly well-suited for medical imaging duties due to its exceptional ability to manage long-range dependencies and sequential information.
@@
State Space Models (SSM) [gu2021combining] describe systems in terms of their internal states and observations over time, thereby facilitating effective sequence modeling through these underlying states. The fundamental form is denoted as follows: $\mathbf{x}_{t}$ is the input state vector, $\mathbf{u}_{t}$ is the control input, $\mathbf{w}_{t}$ is the process noise, $\mathbf{A}$ is the state transition matrix, and $\mathbf{B}$ is the control input matrix.
@@
For observation $\mathbf{y}_{t}$, calculated using the observation matrix $\mathbf{C}$, feedthrough matrix $\mathbf{D}$, and observation noise $\mathbf{v}_{t}$,
@@
The integration of S6 into the Mamba block is especially crucial for sequential medical image processing tasks, such as cardiac MRI segmentation, which require capture of temporal dynamics and structure [gu2023mamba]. The method, on the other hand, is exclusively concerned with per-image segmentation, which involves the application of the state transition and observation matrices ($\mathbf{A}$, $\mathbf{C}$, etc.) to individual images. Each image is treated independently.
U-Mamba capitalizes on Mamba’s linear scaling advantage to improve CNNs’ capacity to simulate long-range dependencies, all while circumventing the high computational costs associated with self-attention mechanisms employed in Transformers [vaswani2017attention] such as ViT [alexey2021] and SwinTransformer [liu2021swin]. The U-Mamba block, which is comprised of two sequential residual blocks followed by a Mamba block, is shown in Figure 2.
@@
An encoder with these blocks is included in the complete U-Mamba network architecture to capture both local features and long-range dependencies, while a decoder composed of residual blocks and transposed convolutions is used to recover detailed local information and resolution. Skip connections are used to connect hierarchical features from the encoder to the decoder. The final decoder output is processed through a $\times1$$\times$1 convolutional layer and a Softmax layer to generate the final segmentation probability map.
III-C Loss Function
Introducing uncertainty into loss functions entails allocating weights to distinct components of the loss according to the estimated uncertainty for each data point [zhao2020uncertainty, zhao2019uncertainty]. This method allows the model to concentrate on learning from more dependable instances while simultaneously reducing the impact of potentially erroneous or ambiguous data. Kendall and Gal introduced the concept of adjusting loss functions by utilizing homoscedastic and heteroscedastic uncertainty [kendall2017uncertainties]. Heteroscedastic uncertainty fluctuates between instances, while homoscedastic uncertainty remains constant across all data points. The model can improve its resilience and precision by focusing on confident predictions and reducing the impact of equivocal ones by adapting its learning process to capitalize on these uncertainties. This optimization enhances overall performance and improves the training process across diverse datasets [hu2023umednerf, wang2024neural, peng2024uncertainty].
@@ III-C Loss Function
    Dice loss [sudre2017]: This region-based metric emphasizes the overlap between predicted and ground truth areas, ensuring accurate preservation of shape and boundary details in segmented regions.
@@
    Cross-Entropy (CE) loss [ma2004]: This distribution-based loss ensures precise categorization of individual pixels, thereby improving classification accuracy.
@@
    Focal loss [lin2017]: This pixel-level loss addresses class imbalance by assigning greater importance to challenging instances, enhancing the model’s ability to manage complex scenarios [hu2020learning, hu2022sum, hu2023rank].
@@
III-D Optimization Strategies for Model Generalization
To improve the U-Mamba model’s generalizaiton in segmenting cardiovascular images, such as those in the ACDC [bernard2018deep], ImageCAS [zeng2023imagecas], and Aorta dataset [imran2024cis, krebs2024volumetric], our methodology employs Sharpness-Aware Minimization (SAM) optimization [foret2021sharpness]. The model’s generalizability is enhanced by the flattening of the loss landscape, implemented by SAM optimization. Despite the fact that conventional optimization techniques are designed to identify the lowest points in the loss landscape, these points are frequently precipitous, which results in inadequate generalization to new data. SAM, on the other hand, identifies gentler minima—regions in the parameter space where the model’s performance remains consistent and is less susceptible to perturbations.
@@ III-D Optimization Strategies for Model Generalization
By applying these two stages iteratively, SAM directs the optimizer toward parameter configurations that are resilient to perturbations, thereby improving generalization and overall performance. The model is able to identify flattened minima in the loss landscape as a result of this approach, which results in improved generalization [foret2021sharpness].
Figure 3: Segmentation results for various methods on sample images from the ACDC dataset [bernard2018deep]. The Mean Squared Error (MSE) between the output segmentation and the ground truth is shown for each method.
Figure 4: Segmentation results for various methods on sample images from the Aorta dataset [imran2024cis, krebs2024volumetric]. The Mean Squared Error (MSE) between the output segmentation and the ground truth is shown for each method.
Figure 5: Segmentation results for various methods on sample images from the ImageCAS dataset [zeng2023imagecas]. The Mean Squared Error (MSE) between the output segmentation and the ground truth is shown for each method.
IV Experiments
IV-A Datasets
IV-A1 The ACDC Dataset
The Automated Cardiac Diagnosis Challenge (ACDC) dataset [bernard2018deep] is a widely recognized benchmark in medical image analysis, particularly for cardiac MRI segmentation. With a total of 300 images and 2,978 slices, this dataset comprises MRI scans from 150 patients, each divided into numerous slices. normal subjects, myocardial infarction, dilated cardiomyopathy, hypertrophic cardiomyopathy, and abnormal right ventricle are the five distinct categories in which the patients are evenly distributed. Each group is distinguished by specific cardiac pathologies.
@@ IV-A1 The ACDC Dataset
IV-A2 The Aorta Dataset
The Aorta dataset [imran2024cis, krebs2024volumetric] is a meticulously annotated compilation of 50 Computed Tomography Angiography (CTA) images that enables the multi-class segmentation of the aorta and its branches. The axial dimensions of these images range from $389\times 389$ pixels to $516\times 516$ pixels, with an average size of $450\times 450$ pixels. The dataset is guaranteed to be consistent in measurement, as each image maintains an isotropic voxel resolution of 1mm x 1mm x 1mm. The average number of axial segments per scan is 695, with image numbers ranging from 578 to 801. This dataset is essential for the development and testing of sophisticated algorithms that are designed to accurately segment the complex vascular structures within the aorta and its branches, thereby establishing a strong foundation for research in medical image analysis and machine learning.
IV-A3 The ImageCAS Dataset
The ImageCAS dataset [zeng2023imagecas] focuses on the segmentation of coronary arteries using CTA images. This dataset contains approximately 1,000 3D CTA images, making it considerably larger than existing public datasets in this domain, which are crucial for diagnosing and assessing coronary artery disease. The dataset is particularly challenging due to the small size and complex branching patterns of the coronary arteries, as well as the motion artifacts introduced by cardiac and respiratory movements. Ground truth annotations include detailed segmentations of the coronary arteries, providing a comprehensive framework for evaluating the performance of segmentation algorithms in detecting and delineating these critical structures.
@@ IV-A3 The ImageCAS Dataset
IV-B Experimental Settings
IV-B1 Evaluation Metrics
We employ the Dice Similarity Coefficient (DSC) as our primary metric to evaluate the performance of segmentation, according to the evaluation protocol described in [maier-hein2022metrics]. The DSC assesses the overlap between the predicted segmentation and the ground truth mask, thereby providing a reliable indication of the model’s accuracy in delineating cardiac structures. The DSC is defined in Eq. (4).
@@ IV-B1 Evaluation Metrics
We also incorporate the Normalized Surface Dice (NSD) metric that is recommended in [maier-hein2022metrics]. The NSD measures the average distance between the predicted segmentation surface and the ground truth surface, normalized by the ground truth surface area. This metric is particularly useful for assessing the spatial accuracy of the segmentation, especially in clinical scenarios where precise boundary delineation is critical. The NSD is defined as follows:
@@ IV-B2 Implementation Details
We conducted the experiments using the PyTorch framework and two NVIDIA A100 Tensor Core GPUs for training. During training, for the ACDC dataset [bernard2018deep], we used a patch size of [20, 256, 224] and a batch size of 4, with the number of pooling operations per axis configured to [2, 5, 5]. In the case of the ImageCAS dataset [zeng2023imagecas], a patch size of [96, 160, 160] and a batch size of 2 were selected, with pooling operations per axis set to [4, 5, 5]. For the Aorta dataset [imran2024cis, krebs2024volumetric], the patch size was [176, 112, 112], also with a batch size of 2, and the pooling operations per axis were set to [4, 4, 4]. The network configuration comprises 6 stages. An initial learning rate of $5\times 10^{-3}$ was utilized, and the training proceeded for 500 epochs. In the SAM optimization, the hyperparameter $\rho$ controlling the perturbation in Eq. (9) was set to 0.05. The focusing parameter $\gamma$ in the focal loss in Eq. (7) was set to 2. The parameter $M$ of the uncertainty-aware loss in Eq. (8) was set to 3, incorporating Diclikee loss, Cross-Entropy loss, and Focal loss.
@@
TABLE I: Performance comparison of our UU-Mamba with leading medical image segmentation methods on the ACDC dataset [bernard2018deep] for the three anatomical regions—the right ventricle (RV), left ventricle (LV), and myocardium (Myo). The evaluation metric is DSC (%).
Method Average RV $\uparrow$ Myo $\uparrow$ LV $\uparrow$ TransUNet [Chen2021TransUNet] 89.71 88.86 84.53 95.73 Swin-Unet [Cao2021SwinUnet] 90.00 88.55 85.62 95.83 nnUNet [Isensee2021nnUNet] 91.61 90.24 89.24 95.36 nnFormer [Zhou2021nnFormer] 92.06 90.94 89.58 95.65 U-Mamba [ma2024umamba] 92.22 91.83 90.22 94.54 Ours 92.79 92.41 90.90 95.04
@@
TABLE II: The U-Mamba backbone was employed to conduct an ablation study of ACDC dataset [bernard2018deep] for the three anatomical regions—the right ventricle (RV), left ventricle (LV), and myocardium (Myo). The study included the following configurations: (1) only the Cross-Entropy (CE) loss, (2) the uncertainty-aware loss without the SAM optimizer, and (3) the proposed UU-Mamba model (uncertainty-aware loss + SAM optimizer). DSC (%) serves as the evaluation metric.
@@
Method Avg. DSC % $\uparrow$ RV DSC % $\uparrow$ Myo DSC % $\uparrow$ LV DSC % $\uparrow$ CE loss 92.263 91.81 90.31 94.67 Uncertainty-aware loss (CE, Dice, Focal) 92.602 92.36 90.51 94.94 Ours 92.787 92.41 90.90 95.04
@@
TABLE III: Ablation study of Aorta dataset [imran2024cis, krebs2024volumetric] on various configurations with U-Mamba backbone: (1) using only the Cross-Entropy (CE) loss, (2) using the uncertainty-aware Loss without the SAM optimizer, and (3) the proposed UU-Mamba model (uncertainty-aware loss + SAM optimizer). The evaluation metric is average DSC (%), NSD (%), and MSE.
Method Avg. DSC % $\uparrow$ Avg. NSD % $\uparrow$ Avg. MSE % $\downarrow$ CE loss 73.761 92.141 0.0960 Uncertainty-aware loss (CE, Dice, Focal) 75.053 91.747 0.0966 Ours 77.084 93.847 0.0906
TABLE IV: Ablation study of ImageCAS dataset [zeng2023imagecas] on various configurations with U-Mamba backbone: (1) using only the Cross-Entropy (CE) loss, (2) using the uncertainty-aware Loss without the SAM optimizer, and (3) the proposed UU-Mamba model (uncertainty-aware loss + SAM optimizer). The evaluation metric is average DSC (%), NSD (%), and MSE.
Method Avg. DSC % $\uparrow$ Avg. NSD % $\uparrow$ Avg. MSE % $\downarrow$ CE loss 79.496 87.490 0.0006784 Uncertainty-aware loss (CE, Dice, Focal) 81.146 88.045 0.0006105 Ours 81.9983 88.771 0.0005903
IV-C Experimental Results
@@
We conduct a comparison of UU-Mamba with five of the state-of-the-art segmentation models—TransUNet [Chen2021TransUNet], Swin-Unet [Cao2021SwinUnet], nnUNet [Isensee2021nnUNet], nnFormer [Zhou2021nnFormer], and U-Mamba [ma2024umamba]—on the ACDC dataset [bernard2018deep]. Transformer-based networks are TransUNet and Swin-Unet, while nnUNet and nnFormer employ CNN-based architectures. U-Mamba is a hybrid architecture that combines components from both Transformer-based and CNN-based networks.
@@
Using the Dice Similarity Coefficient (DSC) as the evaluation metric, we conducted a quantitative evaluation of UU-Mamba against these five 3D heart segmentation models on the ACDC dataset [bernard2018deep]. The segmentation results for the compared algorithms are depicted in Figure 3 on a few images from the ACDC dataset. Table I provides the average DSC scores in all regions, as well as the DSC scores for each model in three cardiac regions: the right ventricle (RV), myocardium (Myo), and left ventricle (LV). Scores demonstrating the greatest performance are indicated in italics.
@@
Our UU-Mamba model surpasses all other methods, attaining the highest overall performance with an average DSC of 92.79%. The DSC scores for each region are as follows: 92.41% for RV, 90.90% for Myo, and 95.04% for LV for each region. The flexibility and efficacy of our model are demonstrated by its exceptional ability to accurately segment the right ventricle and myocardium. Despite a slight decrease in DSC for the left ventricle compared to other models, this is counterbalanced by the highest overall average DSC and the superior scores in other regions, as illustrated in Table I.
@@
In contrast, TransUNet achieves an average DSC of 89.71%, with a relatively lower score for Myo. The average DSC of Swin-Unet is 90.00%, with the maximum DSC for LV and lower performance for RV. The nnUNet model achieves an average DSC of 91.61%, with significant enhancements in Myo segmentation. Providing robust performance in all regions, particularly the myocardium, the nnFormer model obtains an average DSC of 92.06%. The average DSC of U-Mamba of 92.22% indicates substantial improvements in the RV and Myo regions.
@@
The superior performance of our UU-Mamba model in comparison to the other existing models is emphasized by this quantitative evaluation. Our method exhibits the potential to improve the accuracy of 3D heart segmentation in medical imaging, as it has the highest average DSC and notably strong segmentation in the right ventricle and myocardium. The effectiveness of our approach is underscored by the enhancements it achieves over models such as U-Mamba, nnFormer, and nnUNet. This approach employs sophisticated loss functions and optimization techniques to capitalize on global and local features.
@@
TABLE V: Ablation study to compare DSC and NSD values for various anatomical zones and arteries using three methods: (1) only the Cross-Entropy (CE) loss, (2) the uncertainty-aware Loss without the SAM optimizer, and (3) the proposed UU-Mamba model. The best values for each region are highlighted in bold. The table demonstrates the effectiveness of the UU-Mamba model in achieving higher segmentation accuracy across most regions.
DSC (%) NSD (%) Labels CE loss Uncertainty- aware loss UUMamba CE loss Uncertainty- aware loss Ours Zone 0 89.317 87.803 87.604 76.749 70.280 72.291 Innominate 75.151 77.830 78.126 81.251 84.259 84.768 Zone 1 67.936 65.432 65.491 86.647 84.379 85.001 Left Common Carotid 78.001 77.547 78.689 92.852 93.539 93.453 Zone 2 75.043 75.514 73.749 91.645 91.627 90.585 Left Subclavian Artery 83.260 83.578 84.725 99.124 98.742 98.612 Zone 3 74.306 73.513 73.157 94.146 92.939 93.038 Zone 4 79.590 84.233 84.701 89.008 91.987 93.001 Zone 5 88.860 89.725 89.780 95.571 96.879 97.271 Zone 6 71.166 71.119 73.552 97.692 94.694 98.381 Celiac Artery 67.016 66.311 68.691 97.270 96.138 99.158 Zone 7 68.116 71.036 74.067 99.404 99.060 99.622 SMA 69.002 70.117 71.200 88.940 87.300 88.905 Zone 8 77.029 79.027 78.831 100.000 99.576 100.000 Right Renal Artery 74.873 72.898 73.501 98.032 96.949 97.015 Left Renal Artery 67.907 70.187 71.890 96.650 98.323 97.640 Zone 9 90.805 90.118 91.008 99.877 99.636 99.637 Right Common Iliac Artery 78.244 79.078 86.141 95.438 96.317 98.783 Left Common Iliac Artery 75.914 79.623 86.457 97.638 96.370 99.864 Right Internal Iliac Artery 59.568 65.540 66.409 85.795 82.728 88.424 Left Internal Iliac Artery 67.034 64.154 67.791 96.124 90.865 99.069 Right External Iliac Artery 59.216 61.344 72.469 77.545 77.270 87.139 Left External Iliac Artery 59.148 70.506 74.895 81.853 90.306 96.825
@@
Figure 6: Comparison of loss landscapes for models with and without SAM (Sharpness-Aware Minimization) optimization across three datasets: ACDC [bernard2018deep], ImageCAS [zeng2023imagecas], and Aorta [imran2024cis, krebs2024volumetric]. Each image represents the 3D loss landscape of a model trained on one of the datasets, illustrating the effect of SAM optimization on the model’s ability to find flatter minima. The models without SAM exhibit sharper, more erratic loss contours, indicating less stable convergence, whereas the models with SAM show smoother, reflecting enhanced generalization capabilities. This visualization highlights the impact of SAM optimization in improving model robustness and training stability across diverse data conditions.
IV-D Ablation Study